OpenStudy (anonymous):

A quality control inspector found that 43 out of every 50 radios produced were not defective. What percent of the radios were defective?

OpenStudy (campbell_st):

so how many were defective..?

OpenStudy (anonymous):

If there were 43/50 not deffective then there were 50-43=7 deffective. Therefore: \[\frac{defective}{total}=\frac{7}{50}=0.14\] 14% of the radios were defective.

sammixboo (sammixboo):

Hello I am \(\huge\bf Sammi\). \(\LARGE\bf\color{red}W\color{orange}E\color{yellow}L\color{lime}C\color{aqua}O\color{magenta}M\color{darkorchid}E\) \(\large\bf\ TO\) \(\LARGE\bf\color{red}O\color{orange}P\color{yellow}E\color{lime}N\color{aqua}S\color{magenta}T\color{darkorchid}U\color{red}D\color{orange}Y\color{yellow}!\) \(\normalsize\bf\ If~you~have~any~questions~just~type~@sammixboo~in~the\)\(\normalsize\bf\ comments~of~your~question,~and~I~will~be~there~as~soon~as\) \(\normalsize\bf\ possible!\)

sammixboo (sammixboo):